# Difference Decomposition Networks for Infrared Small Target Detection

### v2

@@ Abstract
Weak target signal response is one of the key challenges in infrared small target detection (IRSTD), leading to missing detections, as some dim targets can be completely overwhelmed by the background. Recent studies have widely adopted difference-based methods, which utilize difference features in spatial and temporal domains. These methods rely on fixed and predetermined coefficients to combine these difference features and prevent the targets from overwhelmed by complex background, but fail to extract the correlation information between difference features and original information. To address this, we propose the Basis Decomposition Module (BDM) that can extract correlation information. Building on BDM, we develop the several variants for IRSTD and name them Difference Decomposition Modules (DDMs), which treat the normalized difference results as the basis features, combine them with the coefficient of these features projected onto the input, and achieve precise detection of small infrared targets without requiring a large number of parameters. Besides target enhancement from difference operations, DDMs also demonstrate robustness to diverse transformations. By integrating DDMs into U-shaped architectures, we develop two networks: the Spatial Difference Network (SD2Net) and the Spatiotemporal Difference Network (STD2Net) for single-frame IRSTD (SIRSTD) and multi-frame IRSTD (MIRSTD), respectively. With DDMs, they demonstrate performance competitive with most established networks on the SIRSTD and MIRSTD tasks. Our codes will be made public at: https://github.com/greekinRoma/IRSTD_HC_Platform.
@@ II-A Single-frame Infrared Small Target Detection
Whereas previous models focused primarily on precision, our approach places greater emphasis on lightweight network design, aiming to strike a balance between computational cost and performance. This enables the single-frame model to remain sufficiently lightweight, creating favorable conditions for incorporating temporal information.
@@ II-B Multi-frame Infrared Small Target Detection
We adopt the result-based extraction framework as our foundation because its decoupling from feature-level fusion makes it inherently lightweight and easy to integrate into existing single-frame models.
@@ II-C Difference-based Methods
All of the studies demonstrated the strong power on highlighting the details and obtaining the motion information. Therefore, our method is justified.
@@ III-A1 The procedure of basis decomposition
Figure 3: Robustness of our methods. (a) Illustration of robustness to data value variations. The representation is constructed based on normalized relative differences between neighboring elements, making it insensitive to uniform shifts or scale changes in the input values. (b) Illustration of rotation invariance. Local patterns under different input directions are aligned to an orientation, resulting in identical weight arrangements for rotated versions of the same pattern.
@@ IV-A3 Evaluation Metrics
IV-B Ablation Study in SIRSTD
IV-B1 Impact of Orthogonality and Normalization
@@
IV-B2 Impact of the Dilation Ratios
@@
IV-B3 Impact of the Number of Channels
@@
IV-B4 Module-wise Metrics Performance
@@
IV-C Comparison with State-of-the-art (SOTA) Methods in SIRSTD

### v3

@@ Abstract
Weak target signal responses is the one of the key challenges in infrared small detection (IRSTD), causing targets to be entirely overwhelmed in background clutter and resulting in missed detections. Although conventional difference-based methods enhance target saliency to overcome the challenge, their reliance on fixed weights limits robust performance across diverse background and target patterns. To address this limitation, the Basis Decomposition Module (BDM) is proposed to adaptively extract correlation information between target responses and background patterns, thereby enabling robust target representations under varying conditions. Building on BDMs, several Difference Decomposition Modules (DDMs) are developed for IRSTD. These modules treat the normalized difference results as basis features and combine them with projection coefficients onto the input feature, thereby strengthening robust target representations. By integrating DDMs into U-shaped architectures, we develop two networks: the Spatial Difference Network (SD2Net) and the Spatiotemporal Difference Network (STD2Net) for single-frame IRSTD and multi-frame IRSTD, respectively. With DDMs, they demonstrate performance competitive with most established networks on the single-frame and multi-frame IRSTD tasks. Our codes will be made public at: https://github.com/greekinRoma/IRSTD_HC_Platform.
@@ II-A Single-frame Infrared Small Target Detection
Most existing methods focus primarily on performance, often overlooking the trade-off between accuracy and computational cost. In this work, we propose a lightweight architecture and backbone with low computational complexity, aiming to mitigate the rapidly growing compute budget required in recent years.
@@ II-B Multi-frame Infrared Small Target Detection
Many methods use different spatial and temporal modules, even though they share the same target. Different modules can focus on varied aspects, so one module might relies on background patters while others depend on target structure, making such fusion unreliable. Different from the above works, we utilize the modules based on the same structures for spatial and temporal features, keeping the target representation consistent.
@@ II-C Difference-based Methods
Prior studies excel at enhancing fine details and capturing motion. However, they are often not robust: with fixed weights, even simple input transformations can change the module response and hurt performance. We address this by using dynamic weights that adapt to each input, improving robustness.
@@ III-A1 The procedure of basis decomposition
Figure 3: Robustness of our methods. (a) Invariant to global affine changes $(y=Kx+T)$, where $x$ and $y$ are the input and output feature values, respectively. (b) Illustration of rotation invariance. Rotation does not change the matching weights for each feature.
@@ IV-A3 Evaluation Metrics
IV-B Property Evaluation
We conduct experiments to investigate the properties of our modules. Based on the above analysis, our method has three key properties: invariance to intensity affine changes, rotation invariance, and small-target enhancement.
IV-B1 small-target enhancement
IV-B2 Invariance to intensity affine changes
IV-B3 rotation invariance
IV-C Ablation Study in SIRSTD
IV-C1 Impact of Orthogonality and Normalization
@@
IV-C2 Impact of the Dilation Ratios
@@
IV-C3 Impact of the Number of Channels
@@
IV-C4 Module-wise Metrics Performance
@@
IV-D Comparison with State-of-the-art (SOTA) Methods in SIRSTD